# Evolution Meets Diffusion: Efficient Neural Architecture Generation

### v1

@@ Abstract
Neural Architecture Search (NAS) has gained widespread attention for its transformative potential in deep learning model design. However, the vast and complex search space of NAS leads to significant computational and time costs. Neural Architecture Generation (NAG) addresses this by reframing NAS as a generation problem, enabling the precise generation of optimal architectures for specific tasks. Despite its promise, mainstream methods like diffusion models face limitations in global search capabilities and are still hindered by high computational and time demands. To overcome these challenges, we propose Evolutionary Diffusion-based Neural Architecture Generation (EDNAG), a novel approach that achieves efficient and network-free architecture generation. EDNAG leverages evolutionary algorithms to simulate the denoising process in diffusion models, using fitness to guide the transition from random Gaussian distributions to optimal architecture distributions. This approach combines the optimization strengths of evolutionary strategies and the generation capabilities of diffusion models, enabling rapid and optimal architecture generation. Extensive experiments demonstrate that EDNAG achieves state-of-the-art (SOTA) performance in architecture optimization, with an improvement in accuracy of up to 10.45%. Furthermore, it eliminates the need for time-consuming training and boosts inference speed by an average of 50×, showcasing its exceptional efficiency.
@@ 1 Introduction
Neural Architecture Search (NAS) Dong et al. (2024); Kang et al. (2023) represents an effective approach for the automated design of deep neural network architectures. However, conventional NAS methods are limited by the significant computational costs associated with exploring vast and intricate search spaces. Recently, the paradigm of Neural Architecture Generation (NAG) Nasir et al. (2024); Lukasik et al. (2022) has revolutionized this optimization challenge by re-framing it from an optimization task into a generative task, thus mitigating the limitations imposed by large search spaces. The NAG framework leverages advanced generation techniques to directly produce optimized neural architectures that are finely tuned for specific downstream applications, offering a more scalable and resource-efficient solution.
@@
Previous architecture generation techniques can be categorized into two primary classes: Variational Auto-Encoders (VAE)-based Lin et al. (2023) and Generative Adversarial Networks (GAN)-based Huang et al. (2023) methods. VAE-based methods train an encoder to map architecture features into latent space and a decoder to generate them from latent space, while GAN-based methods employ a generator to produce architectures with the reward from a discriminator. However, both VAE-based and GAN-based methods require the simultaneous training of two networks, leading to significant challenges in high computational complexity, training stability Brock (2018), and generation quality. Recently, diffusion-based methods have demonstrated superior performance by training a single network and generating architectures with multiple denoising steps, thereby improving stability and generation quality. For example, DiffusionNAG An et al. (2023) and DiNAS Asthana et al. (2024) employ a graph diffusion framework to generate architectures through iterative denoising steps, where each denoising iteration requires network inference to predict the noise to be eliminated. However, the generation process is inefficient due to the frequent network inferences and the multi-step denoising process, significantly slowing down the overall generation process. Moreover, when confronted with intricate and expansive search spaces, diffusion models often exhibit limitations in generating neural architectures with global optimality, as shown in Figure 2. Their naive diffusion models frequently converge to local optima, leading to the generation of suboptimal architectures. Therefore, diffusion-based methods still encounter substantial computational complexity during the training phase and exhibit suboptimality in the generation of neural architectures.
Figure 2: Each arrow in the figure represents a denoising step in diffusion models. DiffusionNAG simply adopts diffusion models, which tend to generate suboptimal architectures. In contrast, EDNAG leverages the global optimization capability of evolutionary algorithms to improve the generation process, making it more likely to produce globally optimal architectures.
@@
To the end, we propose a novel and efficient Evolutionary Diffusion-based Neural Architecture Generation Framework, called EDNAG. The core of EDNAG lies in improving efficiency while improving generation quality. Specifically, we introduce evolutionary algorithms to simulate the denoising process in diffusion models. Furthermore, we propose a Fitness-guided Denoising (FD) strategy to produces a new generation of architectures from the previous ones in each denoising iteration. Each newly produced architecture can be conceptualized as a weighted summation of prior samples, with superior architectures assigned higher weights to exert a more prominent influence. Therefore, architecture samples will progressively evolve towards high-fitness subspaces through multiple denoising iterations with FD strategy, thereby generating high-performing and global-optimal neural architectures. In particular, EDNAG performs task-specific neural architecture generation with a neural predictor, serving as a dataset-aware fitness evaluator, to guide the conditional generation process with the FD strategy. Notably, the FD strategy eliminates the dependency on the score network in traditional diffusion models during the denoising process and achieves network-free neural architecture generation, therefore greatly improving generation efficiency.
@@
We performed extensive experiments within NAS-Bench-101 Ying et al. (2019) on CIFAR10 and NAS-Bench-201 Dong and Yang (2020) across four task datasets. These experiments demonstrate that EDNAG not only achieves state-of-the-art (SOTA) performance in architecture generation with up to 10.45% improvement in accuracy, but also exhibits an average 50× speedup compared to previous baselines. Furthermore, we expanded the search space to TransNASBench-101 Duan et al. (2021) and MobileNetV3 Howard et al. (2019). The former search space involves seven distinct downstream computer vision tasks, while the latter features an extremely large search space. Experiments demonstrate that EDNAG consistently achieves state-of-the-art performance in architecture generation, highlighting its adaptability to larger search spaces and its transferability to diverse downstream tasks. Our code is available at https://github.com/cloud-zhoubingye/EvoDiff.git.
In summary, our contributions are as follows:
@@
  • We propose a novel Evolutionary Diffusion-based framework for Neural Architecture Generation, called EDNAG, which simulates the denoising process in diffusion models within an evolutionary paradigm. EDNAG combines the superior generation quality of diffusion models with the conditional guidance capabilities of evolutionary algorithms.
@@
  • We introduce a Fitness-guided Denoising (FD) strategy, which utilizes fitness instead of trained networks to progressively generate offspring architecture samples from previous ones, achieving the first training-free conditional neural architecture generation approach.
@@ 2.1 Neural Architecture Generation
Neural Architecture Generation (NAG) has evolved from VAE-based to GAN-based methods and has recently adopted diffusion-based techniques. NAVIGATOR-D3 Hemmi et al. (2024) is a VAE-based NAG method that extracts the graph features of architectures by an encoder and then reconstruct them through a decoder to generate optimal neural architectures. Following the paradigm of GAN, GA-NAS Such et al. (2020) trains a discriminator to distinguish winning architectures and a generator to produce architectures based on the rewards provided by the discriminator, constituting an adversarial optimization problem. DiffusionNAG represents the groundbreaking application of diffusion models in NAG, producing high-performing architectures through a denoising process with a trained score network. Rather than continuous graph diffusion framework as in DiffusionNAG, DiNAS leverages a discrete graph diffusion model with a score network to generate neural architectures. However, both DiffusionNAG and DiNAS usionNAG and DiNAS require numerous denoising iterations to generate architecture samples, along with frequent network inferences during iterations, leading to inefficiency in the generation process. Their iterative denoising process also suffers from slow convergence and a tendency to stagnate in local optima. Moreover, the aforementioned methods require the training process of one or more networks for generation tasks, which consume significant computational resources, while our method, EDNAG, uniquely achieves a network-free and training-free generation approach.
@@ 2.2 Evolutionary Algorithms
Evolutionary algorithms, designed for optimization problems, serve as population-based heuristic paradigms simulating the natural evolution of species, which are widely appreciated for their global optimization capabilities. Recent studies have revealed potential connections between generative models and evolutionary algorithms, and have begun preliminary attempts to address challenging optimization problems with diffusion models. Zhang et al. (2024) mathematically demonstrates the equivalence between diffusion and evolution, conceptualizing the diffusion process as reversed evolution and the denoising process as evolution. Hartl et al. (2024) improves evolutionary algorithms by integrating diffusion models, enabling the retention of historical information across generations and significantly boosting convergence efficiency. Indeed, the aforementioned studies have largely focused on leveraging the properties of diffusion models to enhance evolutionary algorithms for optimization tasks. However, no prior work has attempted to integrate the principles of evolutionary algorithms into diffusion models for conditional generation tasks. Meanwhile, neural architecture generation (NAG), as a conditional generation task aimed at producing optimal neural architectures, naturally highlights the potential of integrating evolutionary algorithms into neural architecture generation, forming a foundational motivation for EDNAG.
@@ 3.1.1 Diffusion Models
Denoising Diffusion Implicit Model (DDIM) Song et al. (2020) is a representative diffusion model. It involves a forward diffusion process, adding Gaussian noise to original data, and a reverse denoising process, where the model learns to generate original data by reversing the diffusion process.
@@
where $\sigma_{t}$ controlls the degree of random perturbations $\epsilon_{t}\sim\mathcal{N}(0,I)$ during the denoising process. Notably, DDIM leverages a neural network $\epsilon_{\theta}$ to predict the perturbed noise $\hat{\epsilon}_{\theta}\left(x_{t},t\right)$ at time step $t$, guiding the denoising process. By iteratively repeating this denoising process in Equation 2, DDIM progressively generates the final samples. Our method, EDNAG, performs architecture generation based on the denoising process of DDIM, but eliminates the necessity for a network to estimate $\hat{\epsilon}_{\theta}\left(x_{t},t\right)$.
@@ 3.1.2 Evolutionary Algorithms
Evolutionary Algorithms (EAs) Coello Coello (2024); Song et al. (2023) are heuristic frameworks known for their superior optimization capabilities, where Genetic Algorithms (GAs) Alhijawi and Awajan (2024); Katoch et al. (2021) serve as a foundational and influential subset. GAs drive the evolution of populations through fitness-oriented natural selection. In each iteration, the generation of new individuals is facilitated by operators such as mutation and crossover, followed by a fitness-based selection process.
@@ 3.2.1 Overview
Figure 3: Fitness-guided denoising process. EDNAG generates neural architectures through iterative denoising process in FD strategy. In each iteration, FD maps the fitness of $x_{t}$ to the probability density of optimal samples $x_{0}$, subsequently uses the DDIM denoising formula to obtain $x_{t-1}$, and finally employs optimization strategies to select the final samples $x_{t-1}$ from $x_{t}$ and denoised $x_{t-1}$.
@@
3:     Map fitness $f(x_{t})$ to probability desity $g[f(x_{t})]$.
@@
In the following sections, we first explain the fitness-guided denoising (FD) strategy. Next, we describe our optimization strategies for selecting architecture samples.
@@ 3.2.2 Denoising Process of EDNAG
With Equation 3, the denoising process can be transformed as Equation 4.
@@
Before diving into the FD strategy, we compare the generation process of EDNAG with that of traditional diffusion-based methods. Traditional diffusion models in Equation 2, adopted by DiifusionNAG and DiNAS, utilizes $\hat{\epsilon}{\theta}\left(x{t},t\right)$—predicted by a trained network—to guide the generation of subsequent samples. In contrast, EDNAG’s generative process, formulated in Equation 4, is driven by $\hat{x}_{0}$, which is estimated through the training-free fitness-based denoising strategy proposed in this work.
@@
By iteratively applying the FD denoising strategy across $x_{t}$, $x_{t-1}$,…, $x_{0}$, which guides the evolution of architecture samples toward higher-fitness subspaces, EDNAG ultimately generates the global-optimal neural architectures without trained networks.
@@ 3.3 Transferable Neural Predictor for Guidance
For task-specific conditional neural architecture generation, fitness evaluation is critical in guiding the denoising process of generation, which is conducted by a transferable neural predictor. Following DiffusionNAG, our neural predictor comprises three components: a set encoder $p(z_{\mathcal{D}}|\mathcal{D})$Lee et al. (2019), a graph encoder $p(z_{\mathcal{G}}|\mathcal{G})$Zhang et al. (2019), and a linear predictor $p(f(x)|z_{\mathcal{D}},z_{\mathcal{G}})$. In DiffusionNAG, the set encoder $p(z_{\mathcal{D}}|\mathcal{D})$ encodes a small subset of images randomly sampled from the task dataset for each prediction. This naive random sampling for each prediction overlooks the uniformity of dataset distribution, resulting in biased and unstable predictions. To address these issues, we first select a coreset Liu et al. (2023); Guo et al. (2022) from the task dataset and feed it into the set encoder $p(z_{\mathcal{D}}|\mathcal{D})$. Specifically, during prediction, each class of images in the task dataset $\mathcal{T}=\left\{\left(x_{t}^{i},y_{t}^{i}\right)\right\}{i=1}^{|\mathcal{T}|}$ undergoes K-means clustering to select centroids of clusters as coreset $\mathcal{D}=\left\{\left(x_{t}^{i},y_{t}^{i}\right)\right\}{i=1}^{|\mathcal{D}|}$, where $|\mathcal{D}|$ represents the number of samples per class and $|\mathcal{D}|\ll|\mathcal{T}|$. Subsequently, the coreset and the neural architecture are encoded by the set encoder $p(z_{\mathcal{D}}|\mathcal{D})$ and the graph encoder $p(z_{\mathcal{G}}|\mathcal{G})$ respectively to the latent representations $z_{\mathcal{D}}$ and $z_{\mathcal{G}}$. Finally, the linear predictor predicts the performance of the neural architecture. The entire process is detailed in Algorithm 2.

### v3

@@ Abstract
Neural Architecture Search (NAS) has gained widespread attention for its transformative potential in deep learning model design. However, the vast and complex search space of NAS leads to significant computational and time costs. Neural Architecture Generation (NAG) addresses this by reframing NAS as a generation problem, enabling the precise generation of optimal architectures for specific tasks. Despite its promise, mainstream methods like diffusion models face limitations in global search capabilities and are still hindered by high computational and time demands. To overcome these challenges, we propose Evolutionary Diffusion-based Neural Architecture Generation (EDNAG), a novel approach that achieves efficient and network-free architecture generation. EDNAG leverages evolutionary algorithms to simulate the denoising process in diffusion models, using fitness to guide the transition from random Gaussian distributions to optimal architecture distributions. This approach combines the optimization strengths of evolutionary strategies and the generation capabilities of diffusion models, enabling rapid and optimal architecture generation. Extensive experiments demonstrate that EDNAG achieves state-of-the-art (SOTA) performance in architecture optimization, with an improvement in accuracy of up to 10.45%. Furthermore, it eliminates the need for time-consuming training and boosts inference speed by an average of 50×, showcasing its exceptional efficiency. Codes are available at https://github.com/cloud-zhoubingye/EDNAG.git.
@@ 1 Introduction
Neural Architecture Search (NAS) Dong et al. (2024); Kang et al. (2023) represents an effective approach for the automated design of deep neural network architectures. However, conventional NAS methods are limited by the significant computational costs associated with exploring vast and intricate search spaces. Recently, the paradigm of Neural Architecture Generation (NAG) Nasir et al. (2024); Lukasik et al. (2022) has revolutionized this optimization challenge by reframing it from an optimization task into a generation task, thus mitigating the limitations imposed by large search spaces. The NAG framework leverages advanced generation techniques to directly produce optimized neural architectures that are finely tuned for specific downstream applications, offering more scalable and resource-efficient solutions.
@@
Previous architecture generation techniques can be categorized into two primary classes: Variational Auto-Encoders (VAE)-based Lin et al. (2023) and Generative Adversarial Networks (GAN)-based Huang et al. (2023) methods. VAE-based methods train an encoder to map architecture features into embedding representations and a decoder to generate them from the latent space, while GAN-based methods employ a generator to produce architectures with the reward from a discriminator. However, both VAE-based and GAN-based methods require the simultaneous training of two networks, leading to significant challenges in high computational complexity, training stability Brock (2018), and generation quality. Recently, diffusion-based methods have demonstrated superior performance by training a single network and generating architectures with multiple denoising steps, thereby improving stability and generation quality. For example, DiffusionNAG An et al. (2023) and DiNAS Asthana et al. (2024) employ a graph diffusion framework to generate architectures through iterative denoising steps, where each denoising iteration requires network inference to predict the noise to be eliminated. However, the generation process is inefficient due to the frequent network inferences and the multi-step denoising process, which typically requires thousands of steps, significantly slowing down the generation process. Moreover, when confronted with complex and expansive search spaces, diffusion models often exhibit limitations in generating local optimal neural architectures. Their naive diffusion models frequently converge to local optima, leading to the generation of suboptimal architectures. Therefore, diffusion-based methods still encounter substantial computational complexity during the training phase and exhibit sub-optimality in the generation of neural architectures.
@@
To the end, we propose a novel and efficient Evolutionary Diffusion-based Neural Architecture Generation Framework, called EDNAG. The core of EDNAG lies in improving efficiency while improving generation quality. Specifically, we introduce evolutionary algorithms to simulate the denoising process in diffusion models. Furthermore, we propose a Fitness-guided Denoising (FD) strategy to produce a new generation of architectures from the previous ones in each denoising iteration. Each newly produced architecture can be conceptualized as a weighted summation of prior samples, with superior architectures assigned higher weights to exert a more prominent influence. Therefore, architecture samples will progressively evolve towards high-fitness subspaces through multiple denoising iterations with FD strategy, thereby generating high-performing and global-optimal neural architectures. In particular, EDNAG performs task-specific neural architecture generation with a neural predictor, serving as a dataset-aware fitness evaluator to guide the conditional generation process within the FD strategy. Notably, the FD strategy eliminates the dependency on the score network in traditional diffusion models during the denoising process and achieves network-free neural architecture generation, therefore greatly improving generation efficiency.
@@
We perform extensive experiments within NAS-Bench-101 Ying et al. (2019), NAS-Bench-201 Dong and Yang (2020), and NAS-Bench-301 Zela et al. (2020), demonstrating that EDNAG not only achieves state-of-the-art (SOTA) performance in architecture generation with up to 10.45% improvement in accuracy, but also exhibits an average 50$\times$ speedup compared to previous baselines. Furthermore, we expand the search space to TransNASBench-101 Duan et al. (2021) and MobileNetV3 Howard et al. (2019). The former involves seven distinct downstream computer vision tasks, while the latter features an extremely large search space. Experiments demonstrate EDNAG’s adaptability to expansive search spaces and its transferability to diverse downstream tasks.
In summary, our contributions are as follows.
@@
  • We propose a novel Evolutionary Diffusion-based framework for Neural Architecture Generation, called EDNAG, which simulates the denoising process in diffusion models within an evolutionary paradigm. EDNAG combines the superior generation quality of diffusion models with the global optimization capabilities of evolutionary algorithms.
@@
  • We introduce a Fitness-guided Denoising (FD) strategy, which utilizes fitness instead of trained networks to progressively generate offspring architecture samples from previous ones, achieving the first network-free neural architecture generation approach.
@@ 2.1 Neural Architecture Generation
Neural Architecture Generation (NAG) has evolved from VAE-based to GAN-based methods and has recently adopted diffusion-based techniques. NAVIGATOR-D3 Hemmi et al. (2024) is a VAE-based NAG method that extracts the graph features of architectures by an encoder and then reconstructs them through a decoder to generate optimal neural architectures. Following the paradigm of GAN, GA-NAS Such et al. (2020) trains a discriminator to distinguish winning architectures and a generator to produce architectures based on the rewards provided by the discriminator, constituting an adversarial optimization problem. DiffusionNAG represents the groundbreaking application of diffusion models in NAG, producing high-performing architectures through a denoising process with a trained score network. Rather than a continuous graph diffusion framework as in DiffusionNAG, DiNAS leverages a discrete graph diffusion model with a score network to generate neural architectures. However, both DiffusionNAG and DiNAS require numerous denoising iterations to generate architecture samples, along with frequent network inferences during iterations, leading to inefficiency in the generation process. Their iterative denoising process also suffers from slow convergence and a tendency to stagnate in local optima. Moreover, the aforementioned methods require the training process of one or more networks for generation tasks, which consume significant computational resources, while our method, EDNAG, uniquely achieves a network-free and training-free generation approach.
@@ 2.2 Evolutionary Algorithms
Evolutionary algorithms, designed for optimization problems, serve as population-based heuristic paradigms simulating the natural evolution of species, which are widely appreciated for their global optimization capabilities. Recent studies have revealed potential connections between generative models and evolutionary algorithms, and have begun preliminary attempts to address challenging optimization problems with diffusion models. Zhang et al. (2024) mathematically demonstrates the equivalence between diffusion and evolution, conceptualizing the diffusion process as reversed evolution and the denoising process as evolution. Hartl et al. (2024) improves evolutionary algorithms by integrating diffusion models, enabling the retention of historical information across generations. Indeed, the aforementioned studies have largely focused on leveraging the properties of diffusion models to enhance evolutionary algorithms for optimization tasks. However, no prior work has attempted to integrate the principles of evolutionary algorithms into diffusion models for conditional generation tasks. Meanwhile, neural architecture generation (NAG), as a conditional generation task aimed at producing optimal neural architectures, naturally highlights the potential of integrating evolutionary algorithms into neural architecture generation, forming a foundational motivation for EDNAG.
@@ 3.1.1 Diffusion Models
Denoising Diffusion Implicit Model (DDIM) Song et al. (2020) is a representative diffusion model. It involves a forward diffusion process, adding Gaussian noise to the original data, and a reverse denoising process, where the model learns to generate the original data by reversing the diffusion process.
@@
where $\sigma_{t}$ controls the degree of random perturbations $\epsilon_{t}\sim\mathcal{N}(0,I)$ during the denoising process. Notably, DDIM leverages a neural network $\epsilon_{\theta}$ to predict perturbed noise $\hat{\epsilon}_{\theta}\left(x_{t},t\right)$ at time step $t$, guiding the denoising process. By iteratively repeating this denoising process in Equation 2, DDIM progressively generates the final samples. Our method, EDNAG, performs architecture generation based on the denoising process of DDIM, but eliminates the need for a network to estimate $\hat{\epsilon}_{\theta}\left(x_{t},t\right)$.
@@ 3.1.2 Evolutionary Algorithms
Evolutionary Algorithms (EAs) Coello Coello (2024) are heuristic frameworks known for their superior optimization capabilities, where Genetic Algorithms (GAs) Alhijawi and Awajan (2024) serve as a foundational and influential subset. GAs drive the evolution of populations through fitness-oriented natural selection. In each iteration, the generation of new individuals is facilitated by operators such as mutation and crossover, followed by a fitness-based selection process.
@@ 3.2.1 Overview
Figure 2: Fitness-guided denoising process. EDNAG generates neural architectures through iterative denoising process with FD strategy. In each iteration, FD strategy maps the fitness of $x_{t}$ to the probability density of optimal samples $x_{0}$, subsequently uses the DDIM denoising formula to obtain $x_{t-1}$, and finally employs optimization strategies to select the final samples $x_{t-1}$ from $x_{t}$ and denoised $x_{t-1}$.
@@
3:     Map fitness $f(x_{t})$ to probability density $g[f(x_{t})]$.
@@
In the following sections, we first explain the fitness-guided denoising (FD) strategy. Next, we describe our optimization strategies for selecting architecture samples. More detailed derivations and explanations are provided in the Appendix.
@@ 3.2.2 Denoising Process of EDNAG
With Equation 3, the denoising process can be transformed into Equation 4.
@@
By iteratively applying the FD denoising strategy across $x_{t}$, $x_{t-1}$,…, $x_{0}$, which guides the evolution of architecture samples toward higher-fitness subspaces, EDNAG ultimately generates global-optimal neural architectures without trained networks.
@@ 3.3 Transferable Neural Predictor for Guidance
For task-specific conditional neural architecture generation, fitness evaluation is critical in guiding the denoising process of generation, which is conducted by a transferable neural predictor. Following DiffusionNAG, our neural predictor comprises three components: a set encoder $p(z_{\mathcal{D}}|\mathcal{D})$Lee et al. (2019), a graph encoder $p(z_{\mathcal{G}}|\mathcal{G})$Zhang et al. (2019), and a linear predictor $p(f(x)|z_{\mathcal{D}},z_{\mathcal{G}})$. In DiffusionNAG, the set encoder $p(z_{\mathcal{D}}|\mathcal{D})$ encodes a small subset of images randomly sampled from the task dataset for each prediction. This naive random sampling for each prediction overlooks the uniformity of dataset distribution, resulting in biased and unstable predictions. To address these issues, we first select a coreset Liu et al. (2023) from the task dataset and feed it into the set encoder $p(z_{\mathcal{D}}|\mathcal{D})$. Specifically, during prediction, each class of images in the task dataset $\mathcal{T}=\left\{\left(x_{t}^{i},y_{t}^{i}\right)\right\}{i=1}^{|\mathcal{T}|}$ undergoes K-means clustering to select centroids of clusters as coreset $\mathcal{D}=\left\{\left(x_{t}^{i},y_{t}^{i}\right)\right\}{i=1}^{|\mathcal{D}|}$, where $|\mathcal{D}|$ represents the number of samples per class and $|\mathcal{D}|\ll|\mathcal{T}|$. Subsequently, the coreset and the neural architecture are encoded by the set encoder $p(z_{\mathcal{D}}|\mathcal{D})$ and the graph encoder $p(z_{\mathcal{G}}|\mathcal{G})$, respectively, into latent representations $z_{\mathcal{D}}$ and $z_{\mathcal{G}}$. Finally, the linear predictor predicts the performance of the neural architecture. The entire process is detailed in Algorithm 2.